Problem 1. Find the solution to the following initial value problems. (a) y – y – 4y + 4y = 0; y(0) = -4, y(0) = -1, y(
Answer #1

to lochtit, ha ichish 61-=cosch t = (3/h its = 10th ... Auniliary equation is 83-8²-48+4=0 solving we get 8=-2, 2, I .: y ha53 9-49+77 _by = 0, yoou=h, ro)=0, y(0)=0 Aunitiary equation is 83_48²+78-620 solving weget 8=2 tt1.414213562 i :: yona g